Freight Padby 30 FOB / Foo shipping point 1/1000 OB 2/10 Accounting numeric field ote X
Answer #1
Amount paid = Invoice Cost - Customer Returns and Allowances-Discount+Freight
a 30400 =32000-1600
b 10394 =(12800-2500)*(1-2%)+300
c 16830 =(21000-4000)*(1-1%)
d 8015 =(9000-1000)*(1-2%)+175
e 76626 =77400*(1-1%)